# Measuring the Temperature of Extremely Hot Shock-Heated Gas in the Major Merger MACS J0717.5++3745 With Relativistic Corrections to the Sunyaev-Zel’dovich Effect

### v1

@@ Abstract
1 Introduction
The intracluster medium (ICM) comprises approximately $15$ per cent of the total mass and approximately $85$ per cent of the baryonic mass of galaxy clusters (Gonzalez et al., 2007; Lin et al., 2012), making the ICM an important and dynamic constituent in these systems. Direct observations of the structure and thermodynamics of the ICM are critical to understanding the evolution of clusters and their relationship to the underlying dark matter distribution over cosmic time. In the most massive objects ($M\simeq 10^{15}$ M), the ICM can be heated to temperatures near $10^{8}$ K ($kT\simeq 10$ keV) through conversion from potential to kinetic energy as matter accretes onto the cluster (Kravtsov & Borgani, 2012). Additionally, major mergers drive shocks that can heat regions of the ICM to much higher temperatures, with $kT_{\rm ICM}\simeq 40$ keV inferred in extreme systems (e.g., Markevitch, 2006; Markevitch & Vikhlinin, 2007).
The ICM is primarily observed using X-rays where the hot gas emits directly via thermal Bremsstrahlung continuum and line emission (e.g., Bîrzan et al., 2004; Vikhlinin et al., 2006; Markevitch & Vikhlinin, 2007; Cavagnolo et al., 2009; Arnaud et al., 2010). A spectral fit of this emission permits inference of the electron temperature, $T_{e}$ (e.g., Sarazin, 1986; Schellenberger et al., 2015). However, X-ray instruments are typically optimized for energy bands below the highest temperatures expected in the ICM, with sensitivty dropping rapidly by $E\simeq 8$ keV for both XMM-Newton (Jansen et al., 2001) and Chandra (Weisskopf et al., 2000). While not traditionally utilized for ICM studies, X-ray facilities such as NuStar do have the capability to observe photon energies up to $50$ keV (Harrison et al., 2013), but challenges related to its broad point spread function (PSF) and small collecting area limit its ability to measure ICM temperatures (e.g., Wik et al., 2014). As a result of these instrumental limitations, regions of particularly hot gas in major mergers are challenging to characterize via X-rays, which is further compounded by Cosmological dimming in high redshift clusters (e.g., Calzadilla et al., 2024). There is strong motivation to develop alternative ways to measure the temperature of the ICM.
The Sunyaev-Zeldovich effect (SZe) offers a probe of ICM thermodynamics that is independent of and complementary to X-ray observations (Mroczkowski et al., 2019; Lee et al., 2022). The SZe is the spectral distortion in the cosmic microwave background (CMB) radiation caused by inverse-Compton scattering of cold $T_{\rm CMB}=2.73$ K photons from energetic ICM electrons (Sunyaev & Zeldovich, 1972; Birkinshaw, 1999). For a stationary plasma, the effect is dependent on the thermal properties of the gas (tSZe), and can be modeled as an intensity shift in the CMB spectrum described by:
@@ 1 Introduction
where $I_{0}=\frac{2(kT_{CMB})^{3}}{(hc)^{2}}$ and $x\equiv\frac{h\nu}{kT_{CMB}}$ (Carlstrom et al., 2002). The term expressing an integral along the line-of-sight gives the Compton $y$ parameter, which indicates the average energy boost to a CMB photon as a proportion of the electron rest mass. The ICM is sufficiently hot that mild relativistic corrections to the classical scattering are required to fully describe the intensity spectrum (Wright, 1979). These corrections can be included in the term $\delta(x;T_{e})$, a modification that is often referred to as the relativistic correction (rSZe, Itoh et al., 1998; Chluba et al., 2012, 2013; Remazeilles & Chluba, 2025; Remazeilles, 2022). A kinematic effect (kSZe) is also present due to coherent bulk velocities of and in the ICM material along the line of sight, and is modeled by:
$\Delta I_{kSZe}=-I_{0}\frac{x^{4}e^{x}}{(e^{x}-1)^{2}}y(1+\delta(x,v_{\text{pec}},T_{e})).$ (2)
@@
The spectra of both the thermal and kinematic effects are plotted in Fig. 1, assuming properties of the ICM typical of massive clusters. The relativistic corrections for a range of ICM temperatures are also shown, which highlights the temperature dependence of the rSZe corrections. Since the classical (non-relativistic) tSZe brightness is sensitive solely to the product $n_{e}T_{e}$ at a given $x$, it alone cannot be used to isolate the eletron temperature. However, the spectrum of the relativistic correction depends on $T_{e}$, and thus offers a way to probe electron temperature independent of X-ray observations (Remazeilles, 2022).
@@
Figure 1: The spectrum of each component of the SZe for a range of $T_{\rm rSZe}$ calculated from SZPack (Chluba et al., 2012, 2013) assuming a Comptonization $y=2.4\times 10^{-4}$ and a peculiar velocity of $800$ km s-1. Dark grey regions denote the observing bands of Bolocam (Glenn et al., 2003) and red points indicate the band centers of the SPIRE-FTS, both of which are utilized to fit the SZe spectrum in our analysis. Also shown as light grey regions are the observing bands of the Herschel-SPIRE photometer (Valtchanov, 2018), which are used to estimate the contaminating signal from thermal dust emission.
@@
The rSZe is an emerging observable tool that can be used to independently assess the extremely high gas temperatures present in some media, with the shock-heated ICM being an example. The typical small change in SZe amplitude induced by the correction and its spectral degeneracy with the tSZe and kSZe below 300 GHz has restricted its use in traditional radio- and mm-wavelength observations. However, combining sensitive mm-wavelength measurements with sub-mm measurements has been demonstrated to be a potentially fruitful way to isolate the rSZe corrections and explore extreme ICM temperatures (Zemcov et al., 2010, 2012; Butler et al., 2022), and recent statistical measurements using Planck have also utilized the rSZe as a probe of ICM temperature (Hurier & Tchernin, 2017; Erler et al., 2018; Remazeilles & Chluba, 2025).
@@
In this work, we describe an analysis wherein observations from the Herschel Space Observatory’s Spectral and Photometric Imaging REceiver Fourier Transform Spectrometer (SPIRE-FTS; Naylor et al. 2014) are used to constrain the rSZe signal from the quadruple-merger MACS J0717.5++3745 (Ma et al., 2009; Sayers et al., 2013; Adam et al., 2017; Breuer et al., 2024). Between 500–900 GHz the fractional contribution of the rSZe correction is maximized (Fig. 1), and its spectral behavior removes degeneracies with the classical $T_{\rm SZe}$ distortion. These measurements of the SZe spectra with a resolving power $\lambda/\Delta\lambda=R\simeq 30$ from a fully spectrometric instrument provide a space to explore feasibility, assess the effects of various systematic errors, and provide new insights on the state of the cluster’s ICM.
@@
This paper is organized as follows. In Section 2 we describe the primary and ancillary data used in our analysis and how instrument systematics are removed from the data. Section 3 is a detailed presentation of the SPIRE-FTS data reduction. In Section 4 we describe the fitting procedure and in section 5 we discuss our results in the context of past and future rSZe measurements.
@@
2 Data
To measure the rSZe in MACS J0717.5++3745, we combine observations from the SPIRE-FTS (Griffin et al., 2010) with a variety of ancillary observations at a wide range of wavelengths. Following reduction of the SPIRE-FTS measurements to data hypercubes, they are combined with Bolocam (Glenn et al., 2003) intensity measurements to determine the spectrum of the SZe (§4). Next, Herschel-SPIRE photometer observations are used to estimate the contaminating signal contributed by dusty infrared galaxies, and cluster member velocity maps are used as a prior on the ICM velocity (Silich et al., 2024). The corrected intensity spectra are then modeled and various ICM and instrumental parameters are constrained, including $T_{\rm rSZe}$. Finally, the X-ray derived temperatures are combined in a way that makes them comparable to the rSZe measurements. Maps summarizing the data sets are shown in Figure 2, and a compilation of fiducial values from each of the data products listed above at each position shown in Figure 2 are provided in Table 2. In the following section we present a summary of each source of data.
2.1 The Herschel SPIRE-FTS
@@
1342252295 LR I MACS J0717.5++3745 400 2012 Oct 6 $12139$ s
1342252296 LR I MACS J0717.5++3745 400 2012 Oct 6 $12139$ s
1342252297 LR SP Dark Sky 400 2012 Oct 6 $3192$ s
1342252903 LR I MACS J0717.5++3745 400 2012 Oct 7 $12139$ s
1342252904 LR I MACS J0717.5++3745 400 2012 Oct 7 $12139$ s
@@
The SPIRE-FTS instrument employed a Mach-Zehnder interferometer (Naylor et al., 2003) as a Fourier transform spectrometer (FTS) with two input ports. The first FTS input was illuminated by the Herschel telescope, while the second FTS input port was coupled to the Spectrometer Calibrator (SCAL), a cryogenic temperature-controlled source designed to load-match emission from the Herschel primary optics (Dohlen et al., 2000). The SPIRE instrument team found that SCAL was not necessary to compensate for the telescope emission during routine observations, but it noticeably increased the photon noise when heated (Swinyard et al., 2010). As a result, SCAL was left to drift with the temperature of the optical bench around $4.5$ K, and for most SPIRE-FTS observations did not play a role. Here we treat the SPIRE-FTS as a low-resolution spectro-photometer sensitive to continuum emission, therefore the SCAL emission will induce a systematic in our data, as discussed further in Section 3.1.
@@
The SPIRE-FTS output focal plane was populated by independent bolometer arrays coupled to the FTS element via a dichroic beamsplitter. The long wavelength array (termed “SLW”) covered 4471018447-1018 GHz and comprised $19$ detectors in a hexagonal close-packed pattern, as described in Griffin et al. (2010). There are known difficulties connecting the continuum brightness measured in the short wavelength array to SLW (Valtchanov et al., 2018), and since the high frequency data does not constrain the SZe emission in a useful way we opted to ignore them in this study.
@@
An internal beam steering mirror near the input of the spectrometer permitted spatial modulation of the telescope beam and is used to “jiggle” the field of view of the focal plane array to produce fully-sampled images. The observations under study were performed in the intermediate sampling mode of the instrument, which comprises four jiggle positions separated by one beam-width (Benielli et al., 2014). The change in optical configuration was also the source of a strong instrumental systematic, as discussed in Section 3.1.
@@
The fundamental SPIRE-FTS data product is an interferogram acquired over an FTS scan with temporal duration set by the desired spectral resolving power. The interferograms are processed through the Herschel Interactive Processing Environment (HIPE; Ott 2010; Fulton et al. 2016). Our observations (shown in Table 1) are taken in intermediate spatial sampling, $R\sim 30$ “low resolution” (LR) spectral mode. Four hundred interferograms were recorded per jiggle position per observation.
@@
The standard SPIRE-FTS observation set included an observation of a “Dark Sky” field (Pearson et al., 2025) near the north ecliptic pole used to monitor the characteristics of the instrument using a consistent, faint astrophysical background illumination (Hopwood et al., 2015; Marchili et al., 2017). All SPIRE-FTS observations have a companion Dark Sky observation taken in the same observing mode. However, the companion Dark Sky observation for the MACS J0717.5++3745 observations presented here has only one jiggle position, meaning there is no analogue Dark Sky measurement for three of the four optical configurations. We describe the impacts of this discrepancy in Section 3.1.
@@
Table 2: Parameter values obtained from the auxiliary data for each of the seven SPIRE-FTS pointing positions shown in Figure 2. The prior on ICM velocity utilized in our fits, $v_{\text{pec}}$ is derived from data in Section 2.3 and described in Section 4. Electron temperatures from XMM-Newton and Chandra are derived from the data described in Sections 2.4.1 and 2.4.2 using the algorithm presented in Section §2.5. The 140 and 270 GHz intensities were obtained from Bolocam as detailed in Section $2.$
2.2 Bolocam

### v2

@@ Abstract
I Introduction
The intracluster medium (ICM) comprises approximately 1515% of the total mass and approximately 8585% of the baryonic mass of galaxy clusters (Gonzalez et al., 2007; Lin et al., 2012), making the ICM an important and dynamic constituent in these systems. Direct observations of the structure and thermodynamics of the ICM are critical to understanding the evolution of clusters and their relationship to the underlying dark matter distribution over cosmic time. In the most massive objects ($M\simeq 10^{15}$ M), the ICM can be heated to temperatures near $10^{8}$ K ($kT\simeq 10$ keV) through conversion from potential to kinetic energy as matter accretes onto the cluster (Kravtsov and Borgani, 2012). Additionally, major mergers drive shocks that can heat regions of the ICM to much higher temperatures, with $kT_{\rm ICM}\simeq 40$ keV inferred in extreme systems (e.g., Markevitch, 2006; Markevitch and Vikhlinin, 2007).
The ICM is primarily observed using X-rays where the hot gas emits directly via thermal Bremsstrahlung continuum and line emission (e.g., Bîrzan et al., 2004; Vikhlinin et al., 2006; Markevitch and Vikhlinin, 2007; Cavagnolo et al., 2009; Arnaud et al., 2010). A spectral fit of this emission permits inference of the electron temperature, $T_{e}$ (e.g., Sarazin, 1986; Schellenberger et al., 2015). However, X-ray instruments are typically optimized for energy bands below the highest temperatures expected in the ICM, with sensitivity dropping rapidly by $E\simeq 8$ keV for both XMM-Newton (Jansen et al., 2001) and Chandra (Weisskopf et al., 2000). While not traditionally utilized for ICM studies, X-ray facilities such as NuStar do have the capability to observe photon energies up to $50$ keV (Harrison et al., 2013), but challenges related to its broad point spread function (PSF) and small collecting area limit its ability to measure ICM temperatures (e.g., Wik et al., 2014). As a result of these instrumental limitations, regions of particularly hot gas in major mergers are challenging to characterize via X-rays, which is further compounded by cosmological dimming in high-redshift clusters (e.g., Calzadilla et al., 2024). There is strong motivation to develop alternative ways to measure the temperature of the ICM.
The Sunyaev-Zeldovich effect (SZe) offers a probe of ICM thermodynamics that is independent of and complementary to X-ray observations (Mroczkowski et al., 2019; Lee et al., 2022). The SZe is the spectral distortion in the cosmic microwave background (CMB) radiation caused by inverse-Compton scattering of cold $T_{\rm CMB}=2.73$ K photons from energetic ICM electrons (Sunyaev and Zeldovich, 1972; Birkinshaw, 1999). For a stationary plasma, the effect is dependent on the thermal properties of the gas (tSZe), and can be modeled as an intensity shift in the CMB spectrum described by:
@@ I Introduction
where $I_{0}=\frac{2(kT_{CMB})^{3}}{(hc)^{2}}$ and $x\equiv\frac{h\nu}{kT_{CMB}}$ (Carlstrom et al., 2002). The term expressing an integral along the line of sight gives the Compton $y$ parameter, which indicates the average energy boost to a CMB photon as a proportion of the electron rest mass. The ICM is sufficiently hot that mild relativistic corrections to the classical scattering are required to fully describe the intensity spectrum (Wright, 1979). These corrections can be included in the term $\delta(x;T_{e})$, a modification that is often referred to as the relativistic correction (rSZe; Itoh et al., 1998; Chluba et al., 2012, 2013; Remazeilles and Chluba, 2025; Remazeilles, 2022). A kinematic effect (kSZe) is also present due to coherent bulk velocities of and in the ICM material along the line of sight, and is modeled by:
$\Delta I_{kSZe}=-I_{0}\beta\tau\frac{x^{4}e^{x}}{(e^{x}-1)^{2}}.$ (2)
@@
Above, $\beta$ is the ratio of the cluster’s peculiar velocity to the speed of light and $\tau$ is the optical depth of electron scattering. The spectra of both the thermal and kinematic effects are plotted in Fig. 1, assuming properties of the ICM typical of massive clusters. The relativistic corrections for a range of ICM temperatures are also shown, which highlights the temperature dependence of the rSZe corrections. Since the classical (nonrelativistic) tSZe brightness is sensitive solely to the product $n_{e}T_{e}$ at a given $x$, it alone cannot be used to isolate the eletron temperature. However, the spectrum of the relativistic correction depends on $T_{e}$, and thus offers a way to probe electron temperature independent of X-ray observations (Remazeilles, 2022).
@@
Figure 1: The spectrum of each component of the SZe for a range of $T_{\rm rSZe}$ calculated from SZPack (Chluba et al., 2012, 2013) assuming a Comptonization $y=2.4\times 10^{-4}$ and a peculiar velocity of $800$ km s-1. Dark gray regions denote the observing bands of Bolocam (Glenn et al., 2003) and red points indicate the band centers of the SPIRE-FTS, both of which are utilized to fit the SZe spectrum in our analysis. Also shown as light gray regions are the observing bands of the Herschel-SPIRE photometer (Valtchanov, 2018), which are used to estimate the contaminating signal from thermal dust emission.
@@
The rSZe is an emerging observable tool that can be used to independently assess the extremely high gas temperatures present in some media, with the shock-heated ICM being an example. The typical small change in SZe amplitude induced by the correction and its spectral degeneracy with the tSZe and kSZe below 300 GHz has restricted its use in traditional radio- and millimeter-wavelength observations. However, combining sensitive millimeter-wavelength measurements with submillimeter measurements has been demonstrated to be a potentially fruitful way to isolate the rSZe corrections and explore extreme ICM temperatures (Zemcov et al., 2010, 2012; Butler et al., 2022), and recent statistical measurements using Planck have also utilized the rSZe as a probe of ICM temperature (Hurier and Tchernin, 2017; Erler et al., 2018; Remazeilles and Chluba, 2025).
The site of a quadruple-merger, MACS J0717.5++3745 is an excellent testbed for measurements of the rSZe. An analysis of this cluster’s morphology (Ma et al., 2009) was undertaken via Chandra X-ray measurements and spectroscopy of the cluster-member galaxies, and found the overall merger geometry to be the following. There are two primary systems in the SE region that are undergoing a major merger oriented in the plane of sky, and contain regions of coincident shock-heated ICM gas. In the NW, another subcluster appears to be infalling along the line of sight, but contains significant cool dense gas, thus indicating that it has not yet reached pericenter passage. Further NW is a second subcluster that is largely stripped of its gas, which is suggestive of an advanced merger stage near the apocenter. This is supported by radio observations of the cluster undertaken with the Jansky Very Large Array (van Weeren et al., 2017), where they find 100100-$300$ kpc radio filaments in the outskirts of the cluster that trace the shock-heated gas and that the northwest subcluster is ram pressure stripped. In further support of the Ma et al. (2009) picture, Sayers et al. (2013) and Adam et al. (2017) have found large kSZe signals that are coincident with the infalling NW system and with inferred velocities consistent with the cluster-member spectroscopy. Throughout these works, evidence for shock-heated gas is consistently found. In particular, Breuer et al. (2025) find strong evidence for shock-heated gas near the cluster center, coincident with the rSZe signal, thus making it one of the brightest rSZe targets.
@@
In this work, we describe an analysis wherein observations from the Herschel Space Observatory’s Spectral and Photometric Imaging REceiver Fourier Transform Spectrometer (SPIRE-FTS; Naylor et al. 2014) are used to constrain the rSZe signal from the quadruple-merger MACS J0717.5++3745. Between $500$ and $900$ GHz the fractional contribution of the rSZe correction is maximized (Fig. 1), and its spectral behavior removes degeneracies with the classical $T_{\rm SZe}$ distortion. These measurements of the SZe spectra with a resolving power $\lambda/\Delta\lambda=R\simeq 30$ from a fully spectrometric instrument provide a space to explore feasibility, assess the effects of various systematic errors, and provide new insights on the state of the cluster’s ICM.
@@
This paper is organized as follows. In Section II we describe the primary and ancillary data used in our analysis and how instrument systematics are removed from the data. Section III is a detailed presentation of the SPIRE-FTS data reduction. In Section IV we describe the fitting procedure, and in section V we discuss our results in the context of past and future rSZe measurements.
@@
II Data
To measure the rSZe in MACS J0717.5++3745, we combine observations from the SPIRE-FTS (Griffin et al., 2010) with a variety of ancillary observations at a wide range of wavelengths. Following reduction of the SPIRE-FTS measurements to data hypercubes, they are combined with Bolocam (Glenn et al., 2003) intensity measurements to determine the spectrum of the SZe (§IV). Next, Herschel-SPIRE photometer observations are used to estimate the contaminating signal contributed by dusty infrared galaxies, and cluster-member velocity maps are used as a prior on the ICM velocity (Silich et al., 2024). The corrected intensity spectra are then modeled, and various ICM and instrumental parameters are constrained, including $T_{\rm rSZe}$. Finally, the X-ray derived temperatures are combined in a way that makes them comparable to the rSZe measurements. Maps summarizing the data sets are shown in Figure 2, and a compilation of fiducial values from each of the data products listed above at each position shown in Figure 2 is provided in Table 2. In the following section, we present a summary of each source of data.
II.1 The Herschel SPIRE-FTS
@@
1342252295 LR I MACS J0717.5++3745 400 2012 Oct 6 $12,139$ s
1342252296 LR I MACS J0717.5++3745 400 2012 Oct 6 $12,139$ s
1342252297 LR SP Dark Sky 400 2012 Oct 6 $3,192$ s
1342252903 LR I MACS J0717.5++3745 400 2012 Oct 7 $12,139$ s
1342252904 LR I MACS J0717.5++3745 400 2012 Oct 7 $12,139$ s
@@
The SPIRE-FTS instrument employed a Mach-Zehnder interferometer (Naylor et al., 2003) as an FTS with two input ports. The first FTS input was illuminated by the Herschel telescope, while the second FTS input port was coupled to the Spectrometer Calibrator (SCAL), a cryogenic temperature-controlled source designed to load-match emission from the Herschel primary optics (Dohlen et al., 2000). The SPIRE instrument team found that SCAL was not necessary to compensate for the telescope emission during routine observations, but it noticeably increased the photon noise when heated (Swinyard et al., 2010). As a result, SCAL was left to drift with the temperature of the optical bench around $4.5$ K, and for most SPIRE-FTS observations did not play a role. Here, we treat the SPIRE-FTS as a low-resolution (LR) spectro-photometer sensitive to continuum emission; therefore, the SCAL emission will induce a systematic in our data, as discussed further in Section III.1.
@@
The SPIRE-FTS output focal plane was populated by independent bolometer arrays coupled to the FTS element via a dichroic beamsplitter. The long wavelength array (termed “SLW”) covered 4471018447-1018 GHz and comprised $19$ detectors in a hexagonal close-packed pattern, as described in Griffin et al. (2010). There are known difficulties connecting the continuum brightness measured in the short-wavelength array to SLW (Valtchanov et al., 2018), and since the high-frequency data (9441568944-1568 GHz) does not constrain the SZe emission in a useful way, we opted to ignore them in this study.
@@
An internal beam steering mirror near the input of the spectrometer permitted spatial modulation of the telescope beam and is used to “jiggle” the field of view of the focal plane array to produce fully-sampled images. The observations under study were performed in the intermediate sampling mode of the instrument, which comprises four jiggle positions separated by one beamwidth (Benielli et al., 2014). The change in optical configuration was also the source of a strong instrumental systematic, as discussed in Section III.1.
@@
The fundamental SPIRE-FTS data product is an interferogram acquired over an FTS scan with temporal duration set by the desired spectral resolving power. The interferograms are processed through the Herschel Interactive Processing Environment (HIPE; Ott 2010; Fulton et al. 2016). Our observations (shown in Table 1) are taken in intermediate spatial sampling, $R\sim 30$, LR. Four hundred interferograms were recorded per jiggle position per observation.
@@
The standard SPIRE-FTS observation set included an observation of a “Dark Sky” field (Pearson et al., 2025) near the north ecliptic pole used to monitor the characteristics of the instrument using a consistent, faint astrophysical background illumination (Hopwood et al., 2015; Marchili et al., 2017). All SPIRE-FTS observations have a companion Dark Sky observation taken in the same observing mode. However, the companion Dark Sky observation for the MACS J0717.5++3745 observations presented here has only one jiggle position, meaning there is no analogue Dark Sky measurement for three of the four optical configurations. We describe the impacts of this discrepancy in Section III.1.
@@
Table 2: Parameter values obtained from the auxiliary data for each of the seven SPIRE-FTS pointing positions shown in Figure 2. The prior on ICM velocity utilized in our fits, $v_{\text{pec}}$ is derived from data in Section II.4 and described in Section IV. Electron temperatures from XMM-Newton and Chandra are derived from the data described in Sections II.5.1 and II.5.2 using the algorithm presented in Section §II.6. The 140 and 270 GHz intensities were obtained from Bolocam as detailed in Section II.3.
II.2 SPIRE Photometer
The Herschel-SPIRE instrument also contains a set of three photometric channels, which are henceforth referred to as the Herschel-SPIRE photometer (SPIRE-P, Griffin et al. (2006)). These channels have band centers of $600$, $856$, and $1200$ GHz, with beam sizes of $36\arcsec$, $25\arcsec$, and $18\arcsec$, respectively. The photodetectors in use are the same kind of bolometers utilized in the SPIRE-FTS and are arranged in three grids of $43$, $88$, and $139$, which are illuminated by the same optics as the SPIRE-FTS, though the two instruments are mutually exclusive, with only one able to observe at a time. Each of these arrays have a subset of $10$ detectors that are coaligned allowing for easy comparison between the three channels. The photometer has a field of view of $4\arcmin\times 8\arcmin$. During observations, a two-dimensional raster scanning strategy is put into service, where the focal plane is rotated $\pm 42.4\degree$ to account for the underfilled detector arrays (Valtchanov, 2018). We use a SPIRE-P observation to constrain the emission of the cosmic infrared background (CIB) and reduce it from our data, which is described in more detail in §III.3.
II.3 Bolocam